# A Novel Diamond-like Carbon based photocathode for PICOSEC Micromegas detector

### v1

@@ 3.2.2 PEs yield
For the prototype, the amplitude distributions of the signal from single PEs and muon beam were fitted using a polya function[16, 17, 15]. The mean amplitude of single PEs and muons were denoted as Qe and Qmu, respectively. The PEs yield is simply estimated by dividing Qmu by Qe. For DLC photocathodes, the results are shown in Table 1. When the thickness of DLC film is 3 $\mathrm{n}\mathrm{m}$, the optimal performance is obtained, for which the yield is approximately 3.7 Npe/$\mu$. By comparing all the photocathodes, it can be seen that as the thickness of the DLC film increases from 3 $\mathrm{n}\mathrm{m}$ to 10 $\mathrm{n}\mathrm{m}$, the PEs yield decreases, which is consistent with the results of QE measurement. Additionally, under the same conditions, an 18 $\mathrm{n}\mathrm{m}$ CsI photocathode was also tested, while the yield was approximately 10 Npe/$\mu$[6].
@@
Thickness of DLC film (nm) Npe/$\mu$ Detection efficiency 3 3.7 97% 5 3.4% 94% 7.5 2.2% 70% 10 1.7% 68%
The more PEs extracted from the photocathode, the higher the detection efficiency will be obtained. As the thickness of the DLC film increases from 3 $\mathrm{n}\mathrm{m}$ to 10 $\mathrm{n}\mathrm{m}$, the detection efficiency of the prototype for charged particles decreases from 97% to 68%, while the detection efficiency approaches 100% when using CsI photocathode. These results indicate that using a 3 $\mathrm{n}\mathrm{m}$ DLC photocathode, the PICOSEC MM prototype can maintain sufficient detection efficiency for charged particles.
3.2.3 Time resolution
The time resolution of the prototype equipped with different types of DLC photocathode was tested. Of these, the 3 $\mathrm{n}\mathrm{m}$ DLC photocathode exhibits the best performance, followed by the 5 $\mathrm{n}\mathrm{m}$, 7.5 $\mathrm{n}\mathrm{m}$ and 10 $\mathrm{n}\mathrm{m}$ photocathodes in descending order of performance, which is consistent with the results of PEs yield and QE. Figure 10(a) and 10(b) show the time resolution of the prototype equipped with 3 $\mathrm{n}\mathrm{m}$ and 10 $\mathrm{n}\mathrm{m}$ DLC photocathodes, respectively. It can be clearly seen that the time resolution improves with higher pre-amplification voltage, which has the same tendency when using CsI photocathode. With the 3 $\mathrm{n}\mathrm{m}$ DLC photocathode, the time resolution reached up to approximately 42 $\mathrm{p}\mathrm{s}$ at the voltage of 550 $\mathrm{V}$ of pre-amplification gap and 300 $\mathrm{V}$ of amplification gap. With the thickness of the DLC film increases, the time resolution deteriorated. With the 10 $\mathrm{n}\mathrm{m}$ DLC photocathode, the best time resolution was only about 76 $\mathrm{p}\mathrm{s}$, even higher voltage on pre-amplification gap was applied.
@@ 3.2.3 Time resolution
As a comparison, the time resolution of the prototype with an 18 $\mathrm{n}\mathrm{m}$ CsI photocathode was also tested at the same voltage. A time resolution of approximately 28 $\mathrm{p}\mathrm{s}$ was obtained when the voltage of pre-amplification gap and amplification gap were 550 $\mathrm{V}$ and 300 $\mathrm{V}$, respectively. Due to the high PEs extraction efficiency of the CsI photocathode, the prototype can achieve better time resolution compared to the DLC photocathode. However, the robustness of the DLC photocathode is a significant advantage compared to CsI photocathode. As a compromise between time resolution and detector’s operation time, DLC photocathode demonstrates the potential to be the most promising candidate. The aging effect of the DLC photocathode was also studied, as well as that of the CsI photocathode, which is shown in the next section.
3.3 Aging test
3.3.1 Setup
At the typical operating point of the PICOSEC MM, the ratio of ions feedback to the position of photocathode is huge, which will cause non-negligible damage to the CsI photocathode. To verify the resistance of the DLC photocathode to ion bombardment, aging test with high intensity laser was carried out. The test setup is shown in figure 11. A pulsed laser (Passat Lat. Compiler-213) was used to provide laser with a wavelength of 213 $\mathrm{n}\mathrm{m}$. The laser beam is split into two beams. One beam line is recorded by a laser energy meter (COHERENT PM10) which is used to monitor the stability of the laser intensity. The other beam line is sent to the prototype for aging study of the photocathodes. The positive voltage is applied to the anode and mesh electrode, while the photocathode is maintained at 0 $\mathrm{V}$ and connected to a picoammeter (Keithley model 6487). The voltage of pre-amplification gap and the amplification gap are set to 420 $\mathrm{V}$ and 300 $\mathrm{V}$, respectively. The prototype maintains a constant gain by keeping the voltage and other factors fixed.
Figure 11: Schematic diagram and picture of aging test using UV laser
In the test, when the prototype is exposed to the UV laser, the photocathode promptly responds by emitting primary PEs, a direct consequence of the photoelectric effect. These PEs are then further ionized and multiplied in the gas gap, while the ions generated during the ionization process feedback to the position of photocathode and bombard the photocathode film. With the continuous irradiation of laser, the current of feedback ions $I(t)$ from the photocathode and the laser intensity $P(t)$ were simultaneously recorded by the picoammeter and the laser energy meter. Since the gas gain remains unchanged, the normalized current $I(t)/P(t)$ can be used as an indicator of the photocathode’s QE performance. Additionally, the number of charges accumulated on the photocathode can be obtained by integrating the current $I(t)$. The visible degeneration area of the CsI photocathode after aging test was investigated under a microscope, and an irradiated area of approximately 0.06 cm2 was obtained. This area differs from the one indicated in the manual, which is a cycle with a diameter of 2.5 $\mathrm{m}\mathrm{m}$, due to the distortion caused by the laser passing through the optics and the window of the prototype.
3.3.2 Results
The test results are shown in figure 12, where the x-axis is the accumulated charge on the photocathode, and the y-axis is the normalized QE performance of the photocathode. The three curves in figure. 11 represent one CsI photocathode and two DLC photocathodes, with the thicknesses of 18 $\mathrm{n}\mathrm{m}$ (black curve), 3 $\mathrm{n}\mathrm{m}$ (blue curve) and 5 $\mathrm{n}\mathrm{m}$ (red curve), respectively. For each curve, the first data point is normalized to 1. As can be seen from the plot, the performance of the CsI photocathode drops rapidly to around 30% after being bombarded by ions of around 10 $\mathrm{m}\mathrm{C}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$. However, the DLC photocathodes, they can maintain stable QE performance after a slight deterioration. When the accumulated ions on the DLC photocathodes reach tens of $\mathrm{m}\mathrm{C}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$, or even exceed 100 $\mathrm{m}\mathrm{C}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$, their performance degradation is less than 20%. Furthermore, the damaged region of the CsI photocathode after aging can be clearly seen under the microscope, which corresponds to the laser irradiation area, but it is very difficult to find marks on the DLC photocathodes when they are bombarded with even more feedback ions. These results indicate that the PICOSEC MM detector with DLC photocathode can operate longer with stable performance than with CsI photocathode.
@@ 4 Conclusion
The DLC photocathode for the PICOSEC MM detector is introduced, and it has shown promising results with its suitable sensitivity and stability. Its performance is optimized through the manufacturing processes, QE measurement and beam test. In the QE measurement, samples with different thickness are studied, and the results indicate that the optimized thickness of DLC photocathode is approximately 3 $\mathrm{n}\mathrm{m}$. Besides, charged particle beam is used to test the PICOSEC MM prototype equipped with DLC photocathode to verify its performance in practical application environment. A time resolution of around 42 $\mathrm{p}\mathrm{s}$ with a detection efficiency of 97% for 150 $\mathrm{G}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{c}$ muons are obtained when the prototype equipped with a 3 $\mathrm{n}\mathrm{m}$ DLC photocathode. The results in beam test agree well with those of the QE measurement. Additionally, the aging effect of ion bombardment was also tested. The two tested DLC photocathodes maintain stable QE performance after a slight decrease, which is less than 20%, after being bombarded with feedback ions of around 100 $\mathrm{m}\mathrm{C}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$. As a comparison, the CsI photocathode has shown rapid damage when only 10 $\mathrm{m}\mathrm{C}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$ is accumulated on the CsI film. In conclusion, the good properties of no hydrolysis, chemical stability and robustness to ion bombardment of the DLC photocathode demonstrate its great application potential as the photocathode for the PICOSEC MM detector.
@@ Acknowledgments
We acknowledge the support of the Program of National Natural Science Foundation of China (grant number 11935014, 12125505, 12075238, and 11605197); the CERN EP R&D Strategic Programme on Technologies for Future Experiments; the RD51 collaboration, in the framework of RD51 common projects; the Cross-Disciplinary Program on Instrumentation and Detection of CEA, the French Alternative Energies and Atomic Energy Commission; the PHENIICS Doctoral School Program of Université Paris-Saclay, France; the COFUND-FP-CERN-2014 program (grant number 665779); the Fundação para a Ciência e a Tecnologia (FCT), Portugal (CERN/FIS-PAR/0005/2021); the Enhanced Eurotalents program (PCOFUND-GA-2013-600382); the US CMS program under DOE contract No. DE-AC02-07CH11359. The authors wish to thank the Hefei Comprehensive National Science Center for their support.

### v5

@@ 5.2 PEs yield
The amplitude distributions of the prototype for single PEs and muon beam are fitted using a Polya function[21, 22, 20]. The mean amplitude of single PEs and muons are denoted as $Q_{e}$ and $Q_{mu}$, respectively. The PEs yield is simply estimated by dividing $Q_{mu}$ by $Q_{e}$. The PEs yield results are shown in Table 1. When the thickness of DLC film is 3 $\mathrm{n}\mathrm{m}$, the optimal performance is obtained, for which the Npe/$\mu$ equals to 3.7. By comparing all the photocathodes, it can be seen that as the thickness of the DLC film increases from 3 $\mathrm{n}\mathrm{m}$ to 10 $\mathrm{n}\mathrm{m}$, the PEs yield decreases, which is consistent with the results of QE measurement. Additionally, under the same conditions, an 18 $\mathrm{n}\mathrm{m}$ CsI photocathode was also tested, while the Npe/$\mu$ was approximately 10[6].
@@
Thickness of DLC film (nm) Npe/$\mu$ Detection efficiency 3 3.7 97% 5 3.4 94% 7.5 2.2 70% 10 1.7 68%
The more PEs extracted from the photocathode, the higher detection efficiency will be obtained. As the thickness of the DLC film increases from 3 $\mathrm{n}\mathrm{m}$ to 10 $\mathrm{n}\mathrm{m}$, the detection efficiency of the prototype for charged particles decreases from 97% to 68%, while the detection efficiency approaches 100% when using CsI photocathode. These results indicate that using a 3 $\mathrm{n}\mathrm{m}$ DLC photocathode, the PICOSEC MM prototype can maintain sufficient detection efficiency for charged particles.
5.3 Time resolution
The time resolution of the prototype equipped with different types of DLC photocathode was tested. Of these, the 3 $\mathrm{n}\mathrm{m}$ DLC photocathode exhibits the best performance, followed by the 5 $\mathrm{n}\mathrm{m}$, 7.5 $\mathrm{n}\mathrm{m}$ and 10 $\mathrm{n}\mathrm{m}$ photocathodes in descending order of performance, which is consistent with the results of PEs yield and QE. Figure 12(a) and 12(b) show the time resolution of the prototype equipped with 3 $\mathrm{n}\mathrm{m}$ and 10 $\mathrm{n}\mathrm{m}$ DLC photocathodes, respectively. It can be clearly seen that the time resolution improves with higher pre-amplification voltage, which has the same tendency when using CsI photocathode. With the 3 $\mathrm{n}\mathrm{m}$ DLC photocathode, the time resolution reaches up to approximately 42 $\mathrm{p}\mathrm{s}$ at the voltage of 550 $\mathrm{V}$ of pre-amplification gap and 300 $\mathrm{V}$ of amplification gap. With the thickness of the DLC film increases, the time resolution deteriorated. With the 10 $\mathrm{n}\mathrm{m}$ DLC photocathode, the best time resolution is only about 76 $\mathrm{p}\mathrm{s}$, even higher voltage on pre-amplification gap was applied.
@@ 6 Conclusion
The DLC photocathode for the PICOSEC MM detector is introduced, and it has shown promising results with its suitable sensitivity and stability. Its performance is optimized through the manufacturing processes, QE measurements and beam tests. In the QE measurements, samples with different thickness were studied, and the results indicate that the optimized thickness of DLC photocathode is approximately 3 $\mathrm{n}\mathrm{m}$. Besides, the aging effect of ion bombardment was also tested. Both tested DLC photocathodes maintain stable QE performance after a slight decrease, which is less than 20%, after being bombarded with feedback ions of around 100 $\mathrm{m}\mathrm{C}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$. As a comparison, the CsI photocathode has shown rapid damage when only 10 $\mathrm{m}\mathrm{C}\mathrm{/}\mathrm{c}\mathrm{m}^{2}$ is accumulated on the CsI film. Additionally, muon beams was used to test the PICOSEC MM prototype equipped with DLC photocathodes to verify its performance in practical application environment. A time resolution of around 42 $\mathrm{p}\mathrm{s}$ with a detection efficiency of 97% for 150 $\mathrm{G}\mathrm{e}\mathrm{V}\mathrm{/}\mathrm{c}$ muons were obtained when the prototype equipped with a 3 $\mathrm{n}\mathrm{m}$ DLC photocathode. In conclusion, the good properties of no hydrolysis, chemical stability and robustness to ion bombardment of the DLC photocathode demonstrate its great application potential as the photocathode for the PICOSEC MM detector.
@@ Acknowledgments
We acknowledge the support of the Program of National Natural Science Foundation of China (grant number 11935014, 12125505, 12075238); the CERN EP R&D Strategic Programme on Technologies for Future Experiments; the RD51 collaboration, in the framework of RD51 common projects; the Cross-Disciplinary Program on Instrumentation and Detection of CEA, the French Alternative Energies and Atomic Energy Commission; the PHENIICS Doctoral School Program of Université Paris-Saclay, France; the COFUND-FP-CERN-2014 program (grant number 665779); the Fundação para a Ciência e a Tecnologia (FCT), Portugal (CERN/FIS-PAR/0005/2021); the Enhanced Eurotalents program (PCOFUND-GA-2013-600382); the US CMS program under DOE contract No. DE-AC02-07CH11359. The authors wish to thank the Hefei Comprehensive National Science Center for their support.